OpenStudy (anonymous):

Find the solution to the equation 64^(x – 4) = 4^(2x)

Parth (parthkohli):

Can I just explain how to do it?

OpenStudy (anonymous):

64=4^3 64^(x – 4) = 4^(2x) 4^3(x-4)=4^2x 3(x-4)=2x 3x-12=2x 3x-2x=12 x=12

Parth (parthkohli):

We have a confusion between Coco and angela

OpenStudy (anonymous):

64=4x4x4 so it is 4^3

OpenStudy (anonymous):

lol i come back and didn even get to put my input! but thanks

Parth (parthkohli):

 So, 6x - 24 = 4x 2x = 24 x = 12